OpenStudy (anonymous):

ten times the square of a non-zero number is equal t fifty times the number. what is the number?

OpenStudy (anonymous):

please help ill give medals and become your fan

OpenStudy (anonymous):

@pizza4dayz

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

@highschoolmom2010

OpenStudy (highschoolmom2010):

@ganeshie8 this would be up your alley

OpenStudy (anonymous):

i did get the response

OpenStudy (anonymous):

didnt

OpenStudy (astrophysics):

Thanks Jim :)) Ok so we have 10 times the square of a non - zero number, that means \[10 \times n^2\] n here represents the non - zero number. The second part of the question states, that the expression above equals 50 times the same number n. So we have \[10 \times n^2 = 50 \times n\]

OpenStudy (astrophysics):

Does that make sense?

OpenStudy (anonymous):

yes thanks

OpenStudy (astrophysics):

Now just solve for n, you know how to do that right?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

its been a while

OpenStudy (astrophysics):

 can you finish it off to solve for n now? It's the same process as I just did.